# Gaussian Scenes: Pose-Free Sparse-View Scene Reconstruction using Depth-Enhanced Diffusion Priors

### v1

@@ 4.2 Implementation Details
We implement our entire framework in PyTorch 2.3.1 and run all experiments on single A5000 or A6000 GPUs. We render images and depth maps in the 400-600 pixel range as this is closest to the output resolution of Stable Diffusion. We finetuned our diffusion model for $10k$ iterations with batch size $16$ and learning rate 1e-4. To enable classifier-free guidance, we randomly dropout conditioning inputs like text, image, confidence map, etc., with probability of $0.05$ during training.
We fit 3D Gaussians to sparse input images and a DUSt3r point cloud and train it for $1k$ iterations as in  Fan et al. (2024) to obtain our initial Gaussian representation $\mathcal{G}$. The classifier-free guidance scales are set to $s_{I}=2.5$ and $s_{T}=7.5$. We use $t_{max}=0.99$ and linearly decrease $t_{min}$ from $0.98$ to $0.70$. We sample novel-view renders and depth maps for $k=20$ DDIM sampling steps. $w_{d}$ is decayed linearly from $1$ to $0.01$. We also linearly decay the weight of $\mathcal{L}_{sample}$ from $1$ to $0.1$ over $10k$ iterations. GScenes reconstructs the complete 3D scene in about $5$ minutes on a single A6000 GPU.
4.3 Comparison Results
Table 1: Quantitative comparison with state-of-the-art sparse-view reconstruction techniques on classical metrics.
@@ 4.3 Comparison Results
FreeNeRF 11.888 12.877 13.680 0.146 0.180 0.197 0.675 0.654 0.638
RegNeRF 12.297 13.209 13.802 0.147 0.170 0.180 0.668 0.656 0.625
DiffusioNeRF 13.134 16.191 16.732 0.167 0.283 0.337 0.680 0.543 0.530
ZeroNVS* 11.902 11.789 11.729 0.142 0.133 0.128 0.710 0.702 0.694
ReconFusion 15.50 16.93 18.19 0.358 0.401 0.432 0.585 0.544 0.511
CAT3D 16.62 17.72 18.67 0.377 0.425 0.460 0.515 0.482 0.460
InstantSplat 13.657 14.733 14.926 0.222 0.265 0.246 0.603 0.540 0.590
COGS 12.267 12.428 13.030 0.173 0.186 0.201 0.616 0.632 0.596
GScenes 14.976 15.972 16.563 0.284 0.295 0.314 0.593 0.517 0.577
Table 2: Quantitative comparison with few-view reconstruction techniques on metrics suited for generative reconstruction.
@@
FreeNeRF 342.484 343.378 334.805 0.252 0.262 0.268 0.376 0.363 0.357
RegNeRF 343.593 335.606 325.228 0.260 0.276 0.257 0.387 0.383 0.371
DiffusioNeRF 323.047 269.995 246.265 0.230 0.166 0.141 0.395 0.325 0.328
ZeroNVS* 356.395 350.362 343.930 0.283 0.294 0.299 0.433 0.427 0.413
InstantSplat 268.526 226.110 242.853 0.191 0.131 0.140 0.294 0.298 0.310
COGS 231.274 274.827 248.220 0.136 0.183 0.152 0.288 0.310 0.286
GScenes 253.659 220.292 210.255 0.176 0.098 0.106 0.284 0.245 0.241
We report qualitative and quantitative comparisons of GScenes against all related baselines in Fig 5 and Tables 1 and  2. Out of the $8$ baselines, only InstantSplat and COGS are pose-free techniques, while the remaining require ground truth poses for both training and evaluation. Note that COGS relies on ground truth camera intrinsics while GScenes and InstantSplat do not. On the classical metrics (Tab 1), we are third behind the 2 SOTA posed reconstruction methods - CAT3D and ReconFusion. Except for FID and KID on the 3-view split, we outperform all related baselines in Table 2. Due to the unavailability of open-source code, evaluating ReconFusion and CAT3D on these measures is unfortunately not possible.
Base w/o conf. map w/o instruction w/o depth Ours Ground Truth
Counter(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Counter(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Treehill(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 6: Ablation Study with the Counter and Treehill scenes of MipNeRF360. Images above are heldout test-view renders obtained after test-time pose alignment (Sec 3.5). All of our proposed conditionings improve the diffusion model’s ability to inpaint missing details and remove artifacts in novel view renders.
4.4 Ablation Studies
In Fig 6, we ablate different components of our diffusion model. We pick the counter and treehill scenes for this experiment. The Base variant only performs image-to-image diffusion in the RGB space and tries to fill in missing details and remove artifacts based on the partial information available from the novel view render. Conditioning with our pixel-aligned confidence measure leads to better detection and elimination of Gaussian artifacts. Foreground and background artifacts in both the render and depth map are not effectively dealt with in the absence of this conditioning input. Performing diffusion only in the RGB space usually produces depth maps with sharp foreground artifacts. Incorporating joint modeling of RGB and depth through our RGBD diffusion model and subsequent depth optimization during scene reconstruction (Sec 3.4) produces smoother depth maps while also restricting floaters and Gaussian artifacts. As we randomly dropout conditioning prompts (instruction, confidence map, etc.) during fine-tuning, the diffusion model is adept at image synthesis even in the absence of the base instruction prompt. However, an intuitive instruction like “Inpaint low-confidence regions in the image and depth map with coherent details and remove all floaters and Gaussian artifacts.” usually guides the model better in localizing empty regions and Gaussian artifacts.
@@ 4.4 Ablation Studies
4.5 Limitations & Future Work
GScenes is a first step towards a generative solution for pose-free sparse-view reconstruction of large complex scenes. However, it is not free of its fair share of limitations. The quality of the final reconstruction depends heavily on the initial relative pose estimation by the DUSt3r pipeline, and even though the poses are further optimized jointly with Gaussian attributes during training, there are still large differences with the ground truth COLMAP poses estimated from dense views as we show in Tab 3. This limits fair comparison with posed reconstruction methods, as even the test-time pose alignment step (Sec 3.5) cannot compensate for the initial errors in the pipeline. Finally, our diffusion model is limited in its 3D outpainting abilities as, unlike a simplified 2-view setting, there is no explicit view conditioning possible in a general pose-free N-view setting. Existing state-of-the-art posed reconstruction techniques like ReconFusion and CAT3D are able to leverage stronger diffusion priors for scene reconstruction as the diffusion models have view conditioning in the form of pixelNeRF embeddings or “raymap” representations. We hope to address some of these limitations as part of our future work.
5 Conclusion
In this work, we present GScenes where we integrate an instruction-following RGBD diffusion model with a pose-free reconstruction pipeline in InstantSplat to reconstruct a $360^{\circ}$ 3D scene from a few uncalibrated 2D images. We also introduce a pixel-aligned confidence measure to further guide the diffusion model in uncertain regions with missing details and artifacts. Our experiments show that GScenes outperforms existing pose-free reconstruction methods in scene reconstruction and performs competitively with state-of-the-art posed sparse-view reconstruction methods.

### v2

@@ 3.2 Implementation Details
GScenes is implemented in PyTorch 2.3.1 on single A5000/A6000 GPUs. Images and depth maps are rendered at 400-600 pixels to align with Stable Diffusion’s resolution. The diffusion model is finetuned for $100k$ iterations (batch size $16$, learning rate 1e-4) with conditioning element dropout probability of $0.05$ for CFG.
Following InstantSplat, we fit 3D Gaussians to sparse inputs and MASt3r point clouds for $1k$ iterations to obtain $\mathcal{G}$. We use classifier-free guidance scales $s_{I}=s_{C}=3.0$ and sample with $k=20$ DDIM steps. We linearly decay $w_{d}$ from $1$ to $0.01$ and $\mathcal{L}_{sample}$ weight from $1$ to $0.1$ over $10k$ iterations. GScenes completes full 3D reconstruction in approximately $5$ minutes on a single A6000 GPU.
3.3 Comparative Results
Table 1: Quantitative comparison with state-of-the-art sparse-view reconstruction techniques on classical metrics. Red, orange, and yellow indicate best, second-best, and third-best performing methods, respectively.
[…]
@@ 3.3 Comparative Results
[…]
COGS 11.814 12.095 12.555 0.183 0.210 0.228 0.619 0.630 0.605
GScenes 13.809 14.507 14.831 0.265 0.275 0.282 0.547 0.530 0.521
[…]
COGS 11.057 12.350 13.300 0.241 0.282 0.318 0.598 0.564 0.539
GScenes 14.751 15.761 16.199 0.340 0.372 0.394 0.410 0.373 0.360
Table 2: Quantitative comparison with few-view reconstruction techniques on metrics suited for generative reconstruction. Red, orange, and yellow indicate best, second-best, and third-best performing methods.
@@
[…]
COGS 251.694 284.829 270.118 0.158 0.183 0.159 0.297 0.313 0.303
GScenes 163.747 156.368 160.919 0.053 0.052 0.055 0.234 0.227 0.230
[…]
COGS 249.530 224.605 207.782 0.198 0.159 0.128 0.308 0.295 0.283
GScenes 107.750 107.329 105.871 0.036 0.046 0.046 0.193 0.190 0.189
We report qualitative and quantitative comparisons of GScenes against all related baselines in Fig 9, 10 and Tables 1 and  2. Additional qualitative results are provided in Fig 14 and 15 (Sec E). Out of the $8$ baselines, only MASt3R + 3DGS and COGS are pose-free techniques, but COGS relies on ground truth camera intrinsics while GScenes and MASt3R + 3DGS do not. In classical metrics (Tab 1), our baseline MASt3R + 3DGS shows better values than GScenes, as these metrics are usually inadequate for evaluating generative NVS techniques. Nevertheless, we include them for completeness and observe that either our method or the baseline is the best pose-free, open-source method. CAT3D and Reconfusion appear to do better in a posed setting, but we cannot confirm these values nor do a fair comparison in a pose-free setting as they are closed-source. For metrics tailored for generative reconstruction (Tab 2), our method comprehensively outperforms the baseline as well as all related methods for both MipNeRF360 and DL3DV-Benchmark datasets. Due to the unavailability of open-source code, evaluating ReconFusion and CAT3D on these measures is unfortunately not possible. However, we do provide a qualitative comparison with both methods in Sec D. Qualitative comparisons with other open-source posed and pose-free methods clearly show that GScenes is far more adept at generating plausible scene content from partial observations.
3.4 Ablation Studies
In Fig 11 and Tab LABEL:tab:ablation (Sec F), we thoroughly ablate different components of our diffusion model. We pick the bonsai and treehill scenes and their 3-view splits for this experiment. The $1^{st}$ out of 8 columns shows the initial novel-view render obtained from the 3D reconstruction pipeline (MASt3r + 3DGS). The Base variant only performs image-to-image diffusion with no other conditioning, and this already provides a strong baseline for inpainting and artifact elimination in novel-view renders. Without CLIP context guidance, the model fails to adhere to the semantics of the input images when inpainting missing details. Without our confidence measure, the model typically fails to differentiate between image structures and artifacts, often producing implausible images despite context and geometry conditioning. Additionally, we train a Cross-Attn variant where context and geometry conditionings are incorporated using cross-attention instead of FiLM modulation layers in the down and middle blocks of the UNet. Our method achieves similar quality and FID/KID/DISTS scores with more than 3x fewer additional trainable parameters. Similar to Tab 1, GScenes performs a little worse on PSNR, SSIM, and LPIPS compared to the baseline, as these metrics penalize the generation of plausible scene details, slightly dissimilar from the original scene. Nevertheless, the qualitative results clearly show the benefits of our proposed context, geometry, and confidence map conditioning.
4 Conclusion
In this work, we present GScenes where we integrate an image-to-image RGBD diffusion model with a pose-free reconstruction pipeline in MASt3R to reconstruct a $360^{\circ}$3D scene from a few uncalibrated 2D images. We introduce context and geometry conditioning through FiLM layers achieving similar performance as a cross-attention variant. We also introduce a pixel-aligned confidence measure to further guide the diffusion model in uncertain regions with missing details and artifacts. Our experiments show that GScenes outperforms existing pose-free reconstruction techniques in scene reconstruction and performs competitively with state-of-the-art posed sparse-view reconstruction methods.
Appendix A Problem Introduction and Discussion
In the absence of poses estimated from dense observations, we instead rely on recent 3D foundational priors (Leroy et al., 2024) for scene initialization from sparse views and encode its estimated cameras for conditioning a Stable Diffusion UNet (Rombach et al., 2022) during both training and evaluation. We also augment the UNet with additional channels for context, 3DGS render, depth maps with Gaussian artifacts, and a pixel-aligned confidence map capturing missing regions and reconstruction artifacts in the RGBD image. During inference, the model predicts a clean, inpainted version of the conditioning artifact image and an aligned depth map. This formulation prevents the requirement of accurate ground truth poses for pose conditioning of a multiview diffusion model. This also alleviates dependency on large-scale 3D datasets which are usually synthetic and low quality compared to real-world scenes.
We present GScenes, an efficient method that uses 3D foundational (Leroy et al., 2024) and RGBD diffusion priors for pose-free sparse-view reconstruction of complex $360^{\circ}$ scenes. Our method leverages stronger priors than simple regularizers while not relying on million-scale multi-view data or huge compute resources to train a 3D-aware diffusion model. We also ablate all conditioning elements in our diffusion model and identify which conditioning features contribute the most towards coherent NVS from sparse views.
Appendix B Related Work
Reconstructing 3D scenes from limited observations requires generative priors or, more specifically, inpainting missing details in unseen regions in 3D and removing artifacts introduced through observing scene areas from a few observations. Our work builds on recent developments in 2D diffusion priors for 3D reconstruction (Paul et al., 2024), where knowledge learned from abundant 2D datasets is lifted to 3D for refining novel views rendered by sparse 3D models. Next, we discuss how our work is related to the current line of research.
Regularization Techniques
Both NeRF and 3DGS rely on hundreds of scene captures for photorealistic novel view synthesis. When the input set becomes sparse, the problem becomes ill-posed, as several simultaneous 3D representations can agree with the training set. Regularization techniques are amongst the earliest techniques to address this limitation. Typical methods leverage depth from Structure-from-Motion (SfM)(Deng et al., 2022; Roessle et al., 2022), monocular estimation(Li et al., 2024; Xiong et al., 2023; Zhu et al., 2023; Chung et al., 2023), or RGB-D sensors (Wang et al., 2023a). DietNeRF (Jain et al., 2021) uses a semantic consistency loss based on CLIP (Radford et al., 2021) features, while FreeNeRF (Yang et al., 2023) regularizes the frequency range of NeRF inputs. RegNeRF (Niemeyer et al., 2022) and DiffusioNeRF (Wynn & Turmukhambetov, 2023) maximize the likelihoods of rendered patches using normalizing flows or diffusion models, respectively. However, such techniques usually fail under extreme sparsity like 3, 6, or 9 input images for a $360^{\circ}$ scene due to weaker priors. Generative priors can be viewed as a stronger form of regularization as they provide extrapolation capabilities for inferring details in unknown parts of a scene.
Generalizable Reconstruction
When only a few or a single view is available, regularization techniques are often insufficient to resolve reconstruction ambiguities. To address this, recent research focuses on training priors for novel view synthesis across multiple scenes. pixelNeRF (Yu et al., 2021) uses pixel-aligned CNN features as conditioning for a shared NeRF MLP, while other approaches (Trevithick & Yang, 2021; Chen et al., 2021; Henzler et al., 2021; Lin et al., 2023b) condition NeRF on 2D or fused 3D features. Further priors have been learned on triplanes (Irshad et al., 2023), voxel grids (Guo et al., 2022), and neural points (Wewer et al., 2023). Leveraging 3D Gaussian Splatting (Kerbl et al., 2023), methods like pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024) achieve state-of-the-art performance in stereo view interpolation. However, these regression-based techniques infer blurry novel views in case of high uncertainty.
Generative Priors for NVS
While diffusion models have been applied directly on 3D representations like triplanes (Shue et al., 2023; Chen et al., 2023a), voxel grids (Müller et al., 2023), or (neural) point clouds (Zhou et al., 2021; Melas-Kyriazi et al., 2023; Schröppel et al., 2024), 3D data is scarce. Given the success of large-scale diffusion models for image synthesis, there is a great research interest in leveraging them as priors for 3D reconstruction and generation. DreamFusion (Poole et al., 2023) and follow-ups (Wang et al., 2023b; Lin et al., 2023a; Chen et al., 2023b; Deng et al., 2023; Tang et al., 2024) employ score distillation sampling (SDS) to iteratively maximize the likelihood of radiance field renderings under a conditional 2D diffusion prior. For sparse-view reconstruction, existing approaches incorporate view-conditioning via epipolar feature transform (Zhou & Tulsiani, 2023), cross-attention to encoded relative poses (Liu et al., 2023; Sargent et al., 2024), or pixelNeRF (Yu et al., 2021) feature renderings (Wu et al., 2024). However, this fine-tuning is expensive and requires large-scale multi-view data, which we circumvent with GScenes.
Pose-Free 3D Reconstruction
For building a generalizable sparse-view reconstruction method, the assumption of camera poses during inference limits applications to real-world scenarios where usually only an uncalibrated set of 2D images are available with no known camera extrinsics or intrinsics. Several recent works (Jiang et al., 2024; Fan et al., 2024) have attempted to solve reconstruction in a pose-free setting by jointly optimizing poses and NeRF or 3D Gaussian parameters during scene optimization. These methods typically outperform previous techniques (Chen & Lee, 2023; Lin et al., 2021; Bian et al., 2023; Fu et al., 2023) where reconstruction is done in two stages - first estimating poses and then optimizing the 3D representation. However, errors in the initial pose estimation harm subsequent scene optimization, resulting in inferior NVS quality. In our work, we use the MASt3R pipeline with 3DGS for predicting 3D Gaussians and camera parameters in a global coordinate system from a set of unposed 2D images.
Our work is most closely related to $Sp^{2}360$ (Paul et al., 2024) where an instruction-following RGB diffusion model is finetuned for the task of rectifying novel views rendered by 3DGS fitted to sparse observations. We extend the problem setting to the more challenging pose-free scenario and jointly model RGB and depth to aid optimization of 3D Gaussians during the distillation phase. Additionally, we introduce CLIP context and pose conditioning through FiLM layers and a pixel-aligned confidence measure for more accurate novel view synthesis.
Appendix C RGBD Dataset Creation
Our fine-tuning setup for the RGBD diffusion model is illustrated in Fig 12.
Figure 12: Training our RGBD diffusion model. Pairs of RGBD clean images and images with artifacts are obtained from 3DGS fitted to sparse and dense observations, respectively, across $1043$ scenes. CLIP features provide semantic scene context, plücker embeddings of source and target cameras provide geometry information, and a confidence map additionally detects empty regions and artifacts in the artifact image. The Stable Diffusion UNet (Rombach et al., 2022) is then fine-tuned with a dataset of $171,461$ samples.
Appendix D Qualitative Comparison with ReconFusion / CAT3D
MASt3R + 3DGS ReconFusion CAT3D Ours Ground Truth
@@ Appendix D Qualitative Comparison with ReconFusion / CAT3D
Flowers(3) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Plant(3) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Bicycle (9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Qualitative comparison of GScenes with ReconFusion and CAT3D (posed techniques). Despite being a pose-free pipeline built with weaker diffusion priors, our method achieves competitive NVS quality with SOTA sparse-view reconstruction techniques. No image available for CAT3D in the last row, hence kept blank.
In Fig 13, we provide an additional qualitative comparison with ReconFusion and CAT3D. Their code, data, models, or training reproduction details are not available, and hence, we can not perform any evaluation across any of the test scenes in their paper. Despite this issue, from the figures in the 2 papers, we pick the relevant test views for the treehill, flowers, bicycle scenes in MipNeRF360, and the plant scene from CO3Dv2 to show how GScenes compares with their reconstruction. We use the same training views as open-sourced in their data splits. Despite being a pose-free pipeline using weaker generative priors, we observe that GScenes compares competitively with both methods.
Appendix E Additional Qualitative Results
Additional qualitative comparisons on MipNeRF360 and DL3DV-benchmark datasets are provided in Fig 14 and  15.
[…]
Room(3) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Room(3) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Additional Qualitative comparisons on the MipNeRF360 dataset.
[…]
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
6 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Additional Qualitative comparisons on the DL3DV benchmark.
Appendix F Ablation Studies
We report the quantitative performance of our method with different variants of the proposed diffusion model in Tab LABEL:tab:ablation.
Table 3: Ablation study. We report average performance across 3-view splits of bonsai and treehill scenes on 6 metrics. Our proposed diffusion model variant with conditioning from clip features, plücker embeddings, and confidence maps achieves identical FID, KID, and DISTS scores with a more parameter-heavy cross-attention variant. Red, orange, and yellow indicate best, second-best, and third-best performing methods, respectively.
Method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$
Base 12.024 0.271 0.574 264.562 0.105 0.272
Base + Conf. 11.529 0.261 0.572 249.204 0.099 0.274
w/o CLIP 12.166 0.312 0.578 177.707 0.059 0.279
w/o Conf. 12.100 0.270 0.572 265.564 0.131 0.271
Cross-Attn 12.740 0.298 0.543 152.961 0.042 0.225
MASt3R + 3DGS 11.641 0.269 0.604 335.064 0.268 0.329
GScenes 12.871 0.297 0.545 154.769 0.039 0.228
Appendix G Limitations & Future Work
@@
GScenes is a first step towards a generative solution for pose-free sparse-view reconstruction of large complex scenes. However, it is not free of its fair share of limitations. The quality of the final reconstruction depends heavily on the initial relative pose estimation by the MASt3r pipeline, and even though the poses are further optimized jointly with Gaussian attributes during training, there are still large differences with the ground truth COLMAP poses estimated from dense views as we show in Tab 4. This limits fair comparison with posed reconstruction methods, as even the test-time pose alignment step (Sec 3.5) cannot compensate for the initial errors in the pipeline. Our diffusion model, much like related methods, is not agnostic to the 3D representation from which novel view renders, depth, and confidence maps are obtained for fine-tuning the diffusion model for 3D-aware sparse-view NVS. Moreover, our diffusion model trained with 3DGS renders with MASt3R initialization would not be able to rectify novel views from a 3DGS representation with SfM initialization due to the slight difference in the distribution of rendered images. To ensure multiview consistency across all synthesized views, we employ view conditioning in the form of plucker embeddings and use a fixed noise latent across all novel views for multistep reconstruction. However, this does not alleviate the multiview consistency issue completely as novel views are synthesized in an autoregressive manner and not simultaneously synthesized like in video diffusion models. We aim to address these limitations of our diffusion model in future work.